# Optimizing Backward Policies in GFlowNets via Trajectory Likelihood Maximization

### v1

@@ 4.2 Bit Sequences
In this section, we consider the bit sequence generation task introduced by Malkin et al. (2022). Following the experimental setup of (Tiapkin et al., 2024), we modify the state and action spaces to create a non-tree DAG structure, similar to the approach introduced in Zhang et al. (2022).
@@
Importantly, for this environment, the uniform backward coincides with maxent, see Proposition 1 of Zhang et al. (2022) and Remark 3 of Theorem 3 of Mohammadpour et al. (2024).
To evaluate the performance, we use the same metrics as in Malkin et al. (2022) and Tiapkin et al. (2024): the number of modes found during training (number of sequences from $\mathcal{M}$ for which a terminal state within a distance of $30$ has been sampled) and Spearman correlation on the test set between $\mathcal{R}(x)$ and an estimate of $P_{\theta}$. Since computing the exact probability of sampling a terminal state is intractable due to a large number of paths leading to it, we use a Monte Carlo estimate following the approach of Zhang et al. (2022); see also Section A.3.
Figure 2 shows the number of modes for different GFlowNet algorithms and backward approaches found over the course of training. We observe that TLM shows a significant improvement for DB and a minor one for MunchausenDQN in comparison to other backward approaches, where in the later case we find all 60 modes. TB and SubTB also find almost all modes, and TLM does not affect the results much. Full plots for modes across varying learning rates are presented in Figure 5 in Appendix. Figure 3 (top) presents Spearman correlation between $\mathcal{R}$ and $P_{\theta}$ on the test set for the same GFlowNet algorithms and varying learning rates. TLM shows better or similar performance to the baselines across all GFlowNet algorithms if the optimal learning rate is chosen. Moreover, for DB and SubTB, TLM shows steady improvement over the baselines for all learning rates.
Figure 2: Top row: Bit Sequences, the number of discovered modes out of total 60 modes for different methods with learning rate $10^{-3}$. Center row: QM9, the number of Tanimoto-separated modes with reward higher or equal to $1.125$ for different methods with learning rate $5\cdot 10^{-4}$. Bottom row: sEH, the number of Tanimoto-separated modes with reward higher or equal to $0.875$ for different methods with learning rate $5\cdot 10^{-4}$. Higher values indicate better performance.
@@ 4.3 Molecule Design, sEH and QM9
Our final experiments are carried out on molecule design tasks of sEH (Bengio et al., 2021) and QM9 (Jain et al., 2023).
@@
In both tasks, the goal is to generate molecular graphs, with reward emphasizing some desirable property. For both problems, we use pre-trained reward proxy neural networks. For the sEH task, the model is trained to predict the binding energy of a molecule to a particular protein target (soluble epoxide hydrolase) (Bengio et al., 2021). For the QM9 task, the proxy is trained on the QM9 dataset (Ramakrishnan et al., 2014) to predict the HOMO-LUMO gap (Zhang et al., 2020).
@@
For the sEH task, we follow the framework proposed by Jin et al. (2020) and generate molecules using a predefined vocabulary of fragments. It is essential to mention that these fragments are explicitly selected for the sEH task to simplify high-quality object generation. The states are represented as trees of fragments. The actions correspond to choosing a new fragment, then choosing an atom to which the fragment will be attached. There is also a special stop action that moves the state to its terminal copy and stops the generation process.
@@
For QM9 task, molecules are generated atom-by-atom and bond-by-bond. Every state is a connected graph, and actions either add a new node and edge or set the attribute on edge. Thus, the graph-building environment is much more expressive than the tree-building environment, but it results in a more complex generation task and can lead to construction of invalid molecules.
@@
We use the same evaluation metrics for both tasks as proposed in previous works (Madan et al., 2023; Tiapkin et al., 2024). We track the number of Tanimoto-separated modes above a certain reward threshold captured over the course of training, and Pearson correlation on the test set between $\log\mathcal{R}(x)$ and $\log P_{\theta}(x)$. For sEH task we use the same test set as in Bengio et al. (2021), and for QM9 we use a subset of the dataset introduced in Ramakrishnan et al. (2014).
Figure 3: Top row: Bit Sequences, Spearman correlation between $\mathcal{R}$ and $P_{\theta}$ on a test set for different methods and varying learning rate $\in\{5\cdot 10^{-4},10^{-3},2\cdot 10^{-3}\}$. Center row: QM9, Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ on the fixed subset of the QM9 dataset (Ramakrishnan et al., 2014) for different methods and varying learning rate $\in\{5\cdot 10^{-5},10^{-4},5\cdot 10^{-4},10^{-3}\}$. Bottom row: sEH, Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ on the test set from Bengio et al. (2021) for different methods and varying learning rate $\in\{5\cdot 10^{-5},10^{-4},5\cdot 10^{-4},10^{-3}\}$. Higher values indicate better performance.
@@
Figure 2 (center and bottom) shows the number of modes for different GFlowNet algorithms and backward approaches found over the course of training. Overall, TLM greatly speeds up mode discovery on QM9 for all GFlowNet algorithms, but shows similar or worse performance when compared to other backward approaches on sEH. However, we note that on sEH no backward approach shows significant improvement over uniform in terms of mode discovery. Full plots for modes across varying learning rates are presented in Figure 6 in Appendix. It is worth noting that on QM9 TLM shows robust improvement over the baselines across most learning rates for all GFlowNet algorithms. Figure 3 (center and bottom) shows Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ estimate measured on the test set for various learning rates. TLM results in better correlations when paired with MunchausenDQN and SubTB, and shows similar results to the baselines when paired with DB and TB.
Our results support the findings of Mohammadpour et al. (2024), where it was also observed that the backward policy optimization enhances GFlowNets on more complex and less structured environments like QM9.
@@ 5 Conclusion
In this paper, we proposed a new method for backward policy optimization that enhances mode exploration and accelerates convergence in complex GFlowNet environments. This approach represents the first principled method for learning a backward policy in soft reinforcement learning-based GFlowNet algorithms, such as SoftDQN and MunchausenDQN.
In line with the findings of Mohammadpour et al. (2024), we hypothesize that backward policy learning algorithms, such as TLM, may be more advantageous in complex environments with less inherent structure, such as QM9. This is supported by the observation that TLM yields relatively small or negligible improvements in the highly structured sEH task, which has been simplified in prior work (e.g., Jin et al., 2020). In contrast, the unstructured and complex nature of QM9 presents greater challenges, where our approach achieves the most significant performance gains. A deeper understanding of methods like TLM or maxent could enhance the effectiveness of GFlowNets in tackling less structured and more challenging problems.
Another promising direction is using backward policy for exploration as proposed in Kim et al. (2023), He et al. (2024b). Indeed, the ability to sample trajectories that start from high-reward terminal states via $\mathcal{P}_{\mathrm{B}}$ provides an opportunity to improve mode exploration. We expect that combining such methods with TLM-like approaches will additionally improve their performance.
@@ A.2 Stability Techniques
Finally, we find it helpful to initialise $\mathcal{P}_{\mathrm{B}}$ to the uniform distribution at the beggining of training, which is done by zero-initialization of the last linear layer weight and bias.
We ablate the impact of the proposed techniques on QM9, where we try to separately turn off each of the three. Results are presented in Figure 4. We observe that using target model and lower learning rate is crucial, whereas disabling uniform initialization increases variance and shows slightly worse results. For this experiment, we choose DB as the base algorithm because TLM overall shows the greatest impact when applied with it compared to TB, SubTB, and MunchausenDQN.
Figure 4: Ablation study of stability techniques on QM9. The number of Tanimoto-separated modes with a reward at least $1.125$ is shown. As a base algorithm we use DB with a learning rate of $5\cdot 10^{-4}$.
@@ A.3 Experimental Details
We utilize PyTorch (Paszke et al., 2019) in our experiments. For hypergrid and bit sequence environments, we base our implementation upon the published code of Tiapkin et al. (2024). For molecule design experiments, our implementations are based on the open source library by Recursion Pharma.111 https://github.com/recursionpharma/gflownet In all our experiments, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ share the same neural network backbone, predicting logits via two separate linear heads. For all experiments, mean and std values are computed over three random seeds.
@@ A.3.1 Hypergrid
All models are parameterized using an MLP with 2 hidden layers and 256 hidden units. We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 16 trajectories. For the TLM backward policy, we use the same initial learning rate of $10^{-3}$ with an exponential scheduler, tuning $\gamma$ from $\{0.999,0.9999\}$. For SubTB, we set $\lambda=0.9$, following Madan et al. (2023). For SoftDQN and MunchausenDQN, we use a prioritized replay buffer (Schaul et al., 2016) and adopt the same hyperparameters as in Tiapkin et al. (2024). We sample 256 transitions from the buffer to compute the loss for these methods. The backward policy target network is updated using soft updates with a parameter of $\tau=0.25$ (Silver et al., 2014). Since the environment is small, we precompute $\log n(s)$ for all states, allowing us to obtain the maxent backward exactly. Hypergrid experiments were performed on CPUs.
The table summarizes the chosen hyperparameters:
@@ A.3.2 Bit Sequences
All models are parameterized as Transformers Vaswani et al. (2017) with 3 hidden layers, 8 attention heads, and a hidden dimension of 64. Each model is trained for 50,000 iterations and a batch size of 16 with Adam optimizer. We provide results for learning rates from $\{5\times 10^{-4},10^{-3},2\times 10^{-3}\}$. For the TLM backward policy, we use the same initial learning rate as for the forward policy and utilize an exponential scheduler, where $\gamma$ is tuned from $\{0.9997,0.9999\}$. For SubTB we use $\lambda=0.9$. For MunchausenDQN we use a prioritized replay buffer (Schaul et al., 2016) and take the same hyperparameters as in Tiapkin et al. (2024). To compute the loss, we sample 256 transitions from the buffer for MunchausenDQN. The backward policy target network uses soft updates with a parameter $\tau=0.25$ (Silver et al., 2014). Bit sequence experiments were performed on NVIDIA V100 GPUs.
@@
The table summarizes the chosen hyperparameters:
@@ A.3.3 Molecules
We set $\mathcal{R}(x)=\exp(-75.0)$ for invalid molecules $x$ in QM9. In the sEH task, rewards are divided by a constant of 8, and the reward exponent is set to 10. For QM9, we subtract the 95th percentile from all rewards, resulting in the majority of rewards being distributed between 0 and 1, with 5% of molecules having a reward greater than 1. We track the number of Tanimoto-separated modes as described in Bengio et al. (2023), using a Tanimoto similarity threshold of 0.7. After normalization, the reward thresholds are $0.875$ and $1.125$ for sEH and QM9, respectively.
We use the graph transformer architecture from Jain et al. (2023) with 8 layers and 256 embedding dimensions for both tasks. Each model is trained for 20,000 iterations using the Adam optimizer. We present results for learning rates in the set $\{5\times 10^{-4},10^{-4},5\times 10^{-4},10^{-3}\}$. For TLM, the learning rate for $\mathcal{P}_{\mathrm{B}}$ is initially set to be 10 times smaller than that for $\mathcal{P}_{\mathrm{F}}$, while other approaches use the same learning rate for $\mathcal{P}_{\mathrm{B}}$ as for $\mathcal{P}_{\mathrm{F}}$. The learning rate for $\mathrm{Z}$ is fixed at $10^{-3}$ for all experiments. We apply exponential schedulers to all learning rates at each training step and refer readers to the hyperparameter table for the exact decay values. The batch sizes are 256 for sEH and 128 for QM9. The backward policy target network is updated using soft updates with a parameter of $\tau=0.05$ (Silver et al., 2014). For SubTB, we set $\lambda=1.0$, following Madan et al. (2023). For MunchausenDQN, we do not use a replay buffer, training the model on-policy while using the same other hyperparameters as in Tiapkin et al. (2024). To learn $\log n(s)$ for the maxent backward policy, we follow the approach from Mohammadpour et al. (2024), implemented in the Recursion Pharma library. All molecule generation experiments were conducted on NVIDIA A100 GPUs.
In line with previous works (Malkin et al., 2022; Madan et al., 2023; Tiapkin et al., 2024), we use $\varepsilon$-uniform exploration with $\varepsilon=0.05$. To account for the bias introduced into the gradient estimate of $\nabla_{\theta}\mathcal{L}_{\mathtt{TLM}}(\theta_{\mathrm{B}}^{t};\tau)$, we linearly anneal $\varepsilon$ to zero over the course of training. This approach is applied to all GFlowNet algorithms and backward methods to ensure a fair comparison.
The table summarizes chosen hyperparameters:

### v2

@@ 4.2 Bit Sequences
In this section, we consider the bit sequence generation task introduced by Malkin et al. (2022). Following the experimental setup of Tiapkin et al. (2024), we modify the state and action spaces to create a non-tree DAG structure, similar to the approach introduced in Zhang et al. (2022).
@@
To evaluate the performance, we use the same metrics as in Malkin et al. (2022) and Tiapkin et al. (2024): the number of modes found during training (number of sequences from $\mathcal{M}$ for which a terminal state within a distance of $30$ has been sampled) and Spearman correlation on the test set between $\mathcal{R}(x)$ and an estimate of $P_{\theta}$. Since computing the exact probability of sampling a terminal state is intractable due to a large number of paths leading to it, we use a Monte Carlo estimate following the approach of Zhang et al. (2022); see also Section A.3. We train all models with various choices of the learning rate, treating it as a hyperparameter, and provide the results depending on its value, similarly to Madan et al. (2023).
Figure 2 shows the number of modes for different GFlowNet algorithms and backward approaches found over the course of training. We observe that TLM shows a significant improvement for DB and a minor one for MunchausenDQN in comparison to other backward approaches, where in the latter case all backward approaches find all 60 modes. TB and SubTB also find almost all modes, and TLM does not affect the results much in comparison to uniform and naive, while outperforming pessimistic in case of SubTB. Full plots for modes across varying learning rates are presented in Figure 5 in Appendix. Figure 3 (top) presents Spearman correlation between $\mathcal{R}$ and $P_{\theta}$ on the test set for the same GFlowNet algorithms and varying learning rates. TLM shows better or similar performance to the baselines across all GFlowNet algorithms if the optimal learning rate is chosen. Moreover, for DB and SubTB, TLM shows steady improvement over the baselines for all learning rates.
Figure 3: Top row: Bit Sequences, Spearman correlation between $\mathcal{R}$ and $P_{\theta}$ on a test set for different methods and varying learning rate $\in\{5\cdot 10^{-4},10^{-3},2\cdot 10^{-3}\}$. Center row: QM9, Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ on the fixed subset of the QM9 dataset (Ramakrishnan et al., 2014) for different methods and varying learning rate $\in\{5\cdot 10^{-5},10^{-4},5\cdot 10^{-4},10^{-3}\}$. Bottom row: sEH, Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ on the test set from Bengio et al. (2021) for different methods and varying learning rate $\in\{5\cdot 10^{-5},10^{-4},5\cdot 10^{-4},10^{-3}\}$.Higher values indicate better performance. We note here that pessimistic backward policy can be very sensitive to the choice of learning rate.
@@ 4.3 Molecule Design, sEH and QM9
Our final experiments are carried out on molecule design tasks of sEH (Bengio et al., 2021) and QM9 (Jain et al., 2023). In both tasks, the goal is to generate molecular graphs, with reward emphasizing some desirable property. For both problems, we use pre-trained reward proxy neural networks. For the sEH task, the model is trained to predict the binding energy of a molecule to a particular protein target (soluble epoxide hydrolase) (Bengio et al., 2021). For the QM9 task, the proxy is trained on the QM9 dataset (Ramakrishnan et al., 2014) to predict the HOMO-LUMO gap (Zhang et al., 2020).
@@
For the sEH task, we follow the framework proposed by Jin et al. (2020) and generate molecules using a predefined vocabulary of fragments. We use the same 72 fragments as in the seminal work of Bengio et al. (2021). It is essential to mention that these fragments are explicitly selected for the sEH task to simplify high-quality object generation. The states are represented as trees of fragments. The actions correspond to choosing a new fragment and then choosing an atom to which the fragment will be attached. There is also a special stop action that moves the state to its terminal copy and stops the generation process.
@@
For the QM9 task, molecules are generated atom-by-atom and bond-by-bond. Every state is a connected graph, and actions either add a new node and edge or set an attribute on an edge. Thus, the graph-building environment is much more expressive than the fragment-based tree-building environment, but it results in a more complex generation task and can lead to construction of invalid molecules. It is worth mentioning that there exists a simpler fragment-based version of this environment in GFlowNet literature (Shen et al., 2023; Chen & Mauch, 2023), while we consider the more complex atom-by-atom setup from Mohammadpour et al. (2024).
@@
We use the same evaluation metrics for both tasks as proposed in previous works (Bengio et al., 2021; Madan et al., 2023; Tiapkin et al., 2024). We track the number of Tanimoto-separated modes above a certain reward threshold captured over the course of training, and Pearson correlation on the test set between $\log\mathcal{R}(x)$ and $\log P_{\theta}(x)$. For sEH task we use the same test set as in Bengio et al. (2021), and for QM9 we use a subset of the dataset introduced in Ramakrishnan et al. (2014). We train all models with various choices of the learning rate, treating it as a hyperparameter, and provide the results depending on its value, similarly to Madan et al. (2023).
@@
Figure 2 (center and bottom) shows the number of modes for different GFlowNet algorithms and backward approaches found over the course of training. TLM speeds up mode discovery on QM9 when utilized alongside DB and DB and performs on par with the best baseline backward approaches for MunchausenDQN and SubTB, but shows similar or worse performance when compared to other backward approaches on sEH. However, we note that on sEH no backward approach shows significant improvement over uniform in terms of mode discovery. Full plots for modes across varying learning rates are presented in Figure 6 in Appendix. It is worth noting that on the QM9 task, TLM shows robust improvement over the baselines in the majority of configurations. Figure 3 (center and bottom) shows Pearson correlation between $\log\mathcal{R}$ and $\log P_{\theta}$ estimate measured on the test set for various learning rates. TLM results in better correlations when paired with SubTB, shows comaprable results with the best baselines when paired with DB and TB, and falls behind pessimistic when paired with MunchausenDQN while outperforming other backward approaches. However, as Figure 3 indicates, pessimistic backward policy can be very sensitive to the choice of the learning rate and the GFlowNet algorithm that is used for forward policy training.
4.4 Discussion
From the plots above, one can see that across all GFlowNet algorithms (forward policy training objectives), TLM generally shows performance that is better or comparable to other backward approaches. The sole exception is the number of discovered modes in the sEH environment, where TLM can fall behind other backward approaches. To explain this shortcoming, we hypothesize that TLM is more beneficial in environments with less structure. This is supported by the major improvements to mode exploration that it obtains on QM9, while sometimes even degrading the same metric on sEH. Indeed, molecules in the sEH task are constructed from the predefined set of blocks, while in the QM9 task, they are created from atoms. This manually predefined set adds a lot of structure into the environment, in which one creates a junction tree from these blocks. In comparison, the process of creating an arbitrary graph of atoms imposes less structure and can even lead to construction of invalid molecules in certain cases. We suppose that the strong methodological bias is a possible reason why it is of little utility to consider non-trivial backward approaches in the sEH task and why the uniform backward approach often has the best or at least comparable performance according to Figure 6. This is exactly the hypothesis that was initially put forward by Mohammadpour et al. (2024), and our results align with it well.
@@ 5 Conclusion
In this work, we propose a new method for backward policy optimization that enhances mode exploration and accelerates convergence in complex GFlowNet environments. TLM represents the first principled method for learning a backward policy in soft reinforcement learning-based GFlowNet algorithms, such as SoftDQN and MunchausenDQN. We provide an extensive experimental evaluation, demonstrating benefits of TLM when it is paired with various forward policy training methods, and analyze its shortcomings, arguing that our results support the hypothesis of Mohammadpour et al. (2024) about benefits of backward policy optimization in environments with less structure.
A promising further work direction is using backward policy for exploration as proposed in Kim et al. (2023) and He et al. (2024b). Indeed, the ability to sample trajectories starting from high-reward terminal states via $\mathcal{P}_{\mathrm{B}}$ provides an opportunity to improve mode exploration. We expect that combining such methods with TLM-like approaches will additionally improve their performance.
Acknowledgments
The authors are grateful to Dmitry Vetrov for valuable discussions and feedback. The paper was prepared within the framework of the HSE University Basic Research Program. This research was supported in part through computational resources of HPC facilities at HSE University (Kostenetskiy et al., 2021). The work of Daniil Tiapkin was supported by the Paris Île-de-France Région in the framework of DIM AI4IDF.
@@ A.2 Stability Techniques
Finally, we find it helpful to initialize $\mathcal{P}_{\mathrm{B}}$ to the uniform distribution at the beginning of training, which is done by zero-initialization of the last linear layer weight and bias.
We ablate the impact of the proposed techniques on QM9, where we try to turn off each of the three separately. Results are presented in Figure 4. We observe that using a target model and a lower learning rate is crucial, whereas disabling uniform initialization increases variance and shows slightly worse results. For this experiment, we choose DB as the base algorithm because TLM overall shows the most significant impact when applied with it compared to TB, SubTB, and MunchausenDQN.
Figure 4: Ablation study of stability techniques on QM9. The number of Tanimoto-separated modes with a reward at least $1.125$ is shown. As a base algorithm, we use DB with a learning rate of $5\cdot 10^{-4}$.
@@ A.3 Experimental Details
We utilize PyTorch (Paszke et al., 2019) in our experiments. For hypergrid and bit sequence environments, we base our implementation upon the published code of Tiapkin et al. (2024). For molecule design experiments, our implementations are based on the open source library by Recursion Pharmaceuticals.111 https://github.com/recursionpharma/gflownet In all our experiments, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ share the same neural network backbone, predicting logits via two separate linear heads. For all experiments, mean and std values are computed over three random seeds.
An important note is that we do one stochastic gradient update over the replay buffer for pessimistic backward policy approach (Jang et al., 2024) at each training iteration to allow for a fair comparison with other methods. Jang et al. (2024) perform 8 stochastic gradient updates for pessimistic $\mathcal{P}_{\mathrm{B}}$ at each training iteration in the experiments in their paper, which results in a discrepancy in the number of gradient steps when comparing the approach to other methods.
@@ A.3.1 Hypergrid
All models are parameterized using an MLP with 2 hidden layers and 256 hidden units. We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 16 trajectories. For SubTB, we set $\lambda=0.9$, following Madan et al. (2023). For SoftDQN and MunchausenDQN, we use a prioritized replay buffer (Schaul et al., 2016) and adopt the same hyperparameters as in Tiapkin et al. (2024). We sample 256 transitions from the buffer to compute the loss for these methods.
For the TLM backward policy, we use the same initial learning rate of $10^{-3}$ with an exponential scheduler, tuning $\gamma$ from $\{0.999,0.9999\}$. The backward policy target network for TLM uses soft updates with a parameter $\tau=0.25$ (Silver et al., 2014). Since the environment is small, we precompute $\log n(s)$ for all states, allowing us to obtain the maxent backward exactly. For pessimistic backward policy we store trajectories from the last 20 training iterations in the replay buffer and sample 16 trajectories to compute the loss.
Hypergrid experiments were performed on CPUs. Table 1 summarizes the chosen hyperparameters.
Table 1: Hyperparameter choices for hypergrid experiments.
@@ A.3.2 Bit Sequences
All models are parameterized as Transformers Vaswani et al. (2017) with 3 hidden layers, 8 attention heads, and a hidden dimension of 64. Each model is trained for 50,000 iterations and a batch size of 16 with Adam optimizer. We provide results for learning rates from $\{5\times 10^{-4},10^{-3},2\times 10^{-3}\}$. For SubTB we use $\lambda=0.9$. For MunchausenDQN, we use a prioritized replay buffer (Schaul et al., 2016) and take the same hyperparameters as in Tiapkin et al. (2024). To compute the loss, we sample 256 transitions from the buffer for MunchausenDQN.
For the TLM backward policy, we use the same initial learning rate as for the forward policy and utilize an exponential scheduler, where $\gamma$ is tuned from $\{0.9997,0.9999\}$. The backward policy target network for TLM uses soft updates with a parameter $\tau=0.1$ (Silver et al., 2014). For pessimistic backward policy we store trajectories from the last 20 training iterations in the replay buffer and sample 16 trajectories to compute the loss.
@@
Each bit sequence experiment was performed on a single NVIDIA V100 GPU. Table 2 summarizes the chosen hyperparameters.
Table 2: Hyperparameter choices for bit sequence experiments.
@@ A.3.3 Molecules
In the sEH task, rewards are divided by a constant of 8, and the reward exponent is set to 10. For QM9, we subtract the 95th percentile from all rewards, resulting in the majority of rewards being distributed between 0 and 1, with 5% of molecules having a reward greater than 1. The reward exponent is also set to 10. We set $\mathcal{R}(x)=\exp(-75.0)$ for invalid molecules $x$ in QM9. We track the number of Tanimoto-separated modes as described in Bengio et al. (2023), using a Tanimoto similarity threshold of 0.7. After normalization, the reward thresholds are $0.875$ and $1.125$ for sEH and QM9, respectively.
We use the graph transformer architecture from Jain et al. (2023) with 8 layers and 256 embedding dimensions for both tasks. Each model is trained for 20,000 iterations using the Adam optimizer. We present results for learning rates in the set $\{5\times 10^{-4},10^{-4},5\times 10^{-4},10^{-3}\}$. The learning rate for $\mathrm{Z}$ is fixed at $10^{-3}$ for all experiments. We apply exponential schedulers to all learning rates at each training step and refer readers to the hyperparameter table for the exact decay values. The batch sizes are 256 for sEH and 128 for QM9. For SubTB, we set $\lambda=1.0$, following Madan et al. (2023). For MunchausenDQN, we do not use a replay buffer, training the model on-policy while using the same other hyperparameters as in Tiapkin et al. (2024).
For TLM, the learning rate for $\mathcal{P}_{\mathrm{B}}$ is initially set to be 10 times smaller than that for $\mathcal{P}_{\mathrm{F}}$, while other approaches use the same learning rate for $\mathcal{P}_{\mathrm{B}}$ as for $\mathcal{P}_{\mathrm{F}}$. The backward policy target network for TLM uses soft updates with a parameter $\tau=0.05$ (Silver et al., 2014). To learn $\log n(s)$ for the maxent backward policy, we follow the approach from Mohammadpour et al. (2024), implemented in the Recursion Pharmaceuticals GFlowNet library. For pessimistic backward policy we store trajectories from the last 20 training iterations in the replay buffer and sample 256 trajectories to compute the loss for sEH and 128 for QM9 (the same as batch size for these environments).
In line with previous works (Malkin et al., 2022; Madan et al., 2023; Tiapkin et al., 2024), we use $\varepsilon$-uniform exploration with $\varepsilon=0.05$. To account for the bias introduced into the gradient estimate of $\nabla_{\theta}\mathcal{L}_{\mathtt{TLM}}(\theta_{\mathrm{B}}^{t};\tau)$, we linearly anneal $\varepsilon$ to zero over the course of training. This annealing approach is applied to all GFlowNet algorithms and backward methods to ensure a fair comparison in terms of the number of discovered modes.
Each molecule generation experiment was conducted on a single NVIDIA A100 GPU. Table 3 summarizes the chosen hyperparameters.
Table 3: Hyperparameter choices for molecule experiments.